OpenStudy (anonymous):

Find the slope of the line passing through (0,-5) and (-4,7)

OpenStudy (yrelhan4):

Slope= (y2-y1)/(x2-x1)= -5-7/0-(-4) = -12/4 = -3

OpenStudy (anonymous):

Find the slope of the line passing through (5, 5) and (-4, 5).

OpenStudy (yrelhan4):

Do the same way as we did this.. slope=y2-y1/x2-x1..

OpenStudy (anonymous):

then

OpenStudy (yrelhan4):

then what?